# MoVieDrive: Multi-Modal Multi-View Urban Scene Video Generation

### v1

@@ Abstract
Video generation has recently shown superiority in urban scene synthesis for autonomous driving. Existing video generation approaches to autonomous driving primarily focus on RGB video generation and lack the ability to support multi-modal video generation. However, multi-modal data, such as depth maps and semantic maps, are crucial for holistic urban scene understanding in autonomous driving. Although it is feasible to use multiple models to generate different modalities, this increases the difficulty of model deployment and does not leverage complementary cues for multi-modal data generation. To address this problem, in this work, we propose a novel multi-modal multi-view video generation approach to autonomous driving. Specifically, we construct a unified diffusion transformer model composed of modal-shared components and modal-specific components. Then, we leverage diverse conditioning inputs to encode controllable scene structure and content cues into the unified diffusion model for multi-modal multi-view video generation. In this way, our approach is capable of generating multi-modal multi-view driving scene videos in a unified framework. Our experiments on the challenging real-world autonomous driving dataset, nuScenes, show that our approach can generate multi-modal multi-view urban scene videos with high fidelity and controllability, surpassing the state-of-the-art methods.
*D. Huang made contributions during an internship at Huawei.
@@ 1 Introduction
Urban scene video generation for autonomous driving has advanced rapidly in recent years. It can be used to generate controllable driving scenes, especially to synthesize long-tail scenarios that cannot be easily collected in real world. This facilitates performance enhancement and reliability evaluation in autonomous driving.
Contemporary video generation methods, such as SVD [2] and CogVideoX [38], have shown promising performance for generating high-quality videos. However, they cannot be directly used for autonomous driving because multi-view driving scene generation requires multi-view spatiotemporal consistency and high controllability. To address this problem, some recent studies have explored diffusion models [10, 24] for controllable multi-view urban scene generation and shown promising performance [29, 33, 6, 21]. However, these methods mostly focus on single-modal RGB video generation and lack the ability to support multi-modal video generation. As a perception-intensive task, autonomous driving has an inherent need for multi-modal data, such as depth maps and semantic maps, which can facilitate holistic urban scene understanding towards safer and more efficient self-driving. A common approach to solving this problem is using multiple models to generate different modalities, but this often increases the difficulty of model deployment and does not fully exploit complementary cues for multi-modal data generation.
Figure 1: An illustration of different approaches to multi-modal multi-view driving scene generation. (a) Previous approaches typically require multiple models to generate different modalities, while (b) our approach employs a unified model for multi-modal multi-view joint generation.
@@
In this work, we propose a novel multi-MOdal multi-VIEw video diffusion approach, dubbed MoVieDrive, to autonomous driving. As illustrated in Fig. 1, instead of using multiple models to generate different modalities, our approach proposes to employ a unified model for multi-modal multi-view driving scene generation. An overview of our approach is depicted in Fig. 2. Specifically, we employ diverse conditioning inputs, including text conditions, reference conditions, and layout conditions, to guide controllable scene generation. Among them, text conditions guide holistic scene generation, reference conditions are optional and only used for future scene prediction, and layout conditions capture fine-grained scene cues. Next, we decompose multi-modal multi-view scene generation into modal-shared learning and modal-specific learning, and devise a unified diffusion transformer model with modal-shared layers and modal-specific layers. This unified model alleviates the need for multiple models and improves model scalability. We then encode diverse conditions into the unified diffusion model for multi-modal multi-view scene generation. This helps to leverage complementary cues from multi-modal data for richer scene understanding to facilitate multi-modal multi-view urban scene generation in a unified framework. To evaluate the performance of our approach, we conduct extensive experiments on nuScenes [3], a challenging real-world autonomous driving dataset. Experimental results show that our approach achieves compelling video generation quality and controllability compared with the state-of-the-art methods, while supporting multi-modal data generation. In summary, our contributions are:
  • We propose the first work that exploits diverse conditioning inputs and a diffusion transformer model for multi-modal multi-view autonomous driving scene generation. This resolves the limitation of existing works and facilitates urban scene understanding in autonomous driving.
  • We devise a unified multi-modal multi-view diffusion transformer model composed of modal-shared components and modal-specific components. This enables multi-modal multi-view urban scene generation in a unified model.
@@ 2 Related Work
The rapid advancement in urban scene and asset generation has significantly facilitated data synthesis and closed-loop evaluation in autonomous driving [37, 41, 7, 18]. There are two primary research directions in urban scene synthesis for autonomous driving. One line of work explores rendering methods, such as neural radiance fields and 3D Gaussian splatting, for urban scene synthesis [37, 41, 23, 34]. Despite high-fidelity scene synthesis, this type of method has poor controllability and diversity. Another line of work explores video generation, such as diffusion models [2, 38], for urban scene generation [33, 29]. Among them, some studies focus on single-view generation [8, 30], while the others investigate multi-view generation [7, 31, 21]. Our work focuses on the more challenging multi-view setting. Diffusion models are widely used as the backbone for contemporary multi-view urban scene generation, such as DriveDreamer [29], Panacea [33], MagicDrive [7], MagicDrive-V2 [6], Drive-WM [31], MaskGWM [21], etc. Although these methods have shown promising results, they only focus on single-modal RGB video generation and lack the ability to generate multi-modal data for holistic scene understanding. In addition, UniScene [13] proposes using multiple models to generate RGB videos and LiDAR point clouds, but still fails to construct a unified model for multi-view multi-modal driving scene generation. Different from existing works, our approach proposes to generate multi-modal multi-view driving scene videos in a unified framework, which fills a gap left by existing methods.
@@
There have been some diffusion-based multi-modal synthesis methods proposed in recent years [17, 39, 26, 35]. However, none of them are designed for urban scene generation and modifying them for complex controllable multi-modal multi-view scene generation is not trivial. Our approach differs from these works in that we propose a novel framework that uses diverse conditioning inputs to encode controllable scene structure and content cues into a diffusion transformer model for multi-modal multi-view driving scene generation.
@@ 3.1 Approach Overview
Fig. 2 depicts an overview of our approach. Our approach follows the paradigm of diffusion-based video generation [6, 29]. As shown in the left of Fig. 2, we construct the scene descriptions with text conditions, reference conditions, and layout conditions, and employ the corresponding encoders to extract the embeddings of these conditions. The embeddings of layout conditions and reference conditions are concatenated with noisy latents $x$ as inputs $z$ to a diffusion model, while the embeddings of text conditions are injected into the diffusion model through cross-attention layers. Here, noisy latents $x$ are obtained by sampling random noises from the Gaussian distribution during inference and by adding scheduled noises to the groundtruth frame latents extracted by a VAE encoder during training. Next, as shown in the middle of Fig. 2, we construct a multi-modal multi-view diffusion transformer model with modal-shared layers, composed of temporal layers and multi-view spatiotemporal blocks, and modal-specific layers, composed of cross-modal interaction layers and projection layers. We then train the diffusion transformer model to estimate the added noises $\epsilon{\sim}\mathcal{N}(0,I)$ from the conditioned inputs $z$, so as to generate clean latents $x^{\prime}$ for each modality. Finally, as shown in the right of Fig. 2, the clean latents $x^{\prime}$ are used as inputs to a VAE decoder to produce multi-view scene videos for each modality. In the following sections, we present the details of conditioning inputs encoding, multi-modal multi-view diffusion transformer, and model training and inference.
@@ 3.2 Conditioning Inputs Encoding
Following [6, 29], text conditions are used to guide holistic scene generation. In our framework, we employ two types of text conditions, including camera prompts and text prompts, and use dual text encoders to generate text condition embeddings $f^{text}$. For camera prompts $c^{cam}$, we concatenate the camera intrinsic and extrinsic parameters applied with the Fourier embedding [19] and employ an MLP-based encoder $E^{cam}$ to extract camera embeddings. For text prompts $c^{text}$, we use video caption descriptions and apply a frozen T-5 text encoder [22] $E^{text}$ to extract text embeddings. We concatenate these embeddings as $f^{text}$. This process can be formulated as:
@@
For fine-grained scene structure and content control, we employ three types of layout conditions, including box maps, road maps and occupancy-based layout maps. These layout conditions can be obtained from human annotations or pretrained models and are useful to control fine-grained details in complex multi-view driving scene generation. For box maps, different from [7, 6] which uses box coordinates to extract box embeddings, we directly generate 2D box maps $c^{b}$ by projecting target 3D box locations onto image planes and assign different colors to different categories. For road maps $c^{r}$, we project road structure cues (lane dividers, pedestrian crossings and lane boundaries) onto image planes and assign different colors to different categories. In addition, following [14], we use occupancy-based layout maps $c^{o}$ to further enhance scene fine-grained control. This is accomplished by projecting sparse 3D occupancy onto an image plane to generate sparse semantic maps. Note that these semantic maps are coarse and sparse, which are different from the dense semantic modality used for urban scene understanding in autonomous driving. 3D occupancy has been well explored in autonomous driving, and in practice, 3D occupancy can be obtained from an off-the-shelf model [13] or a simulator [5].
@@
where $E^{l}_{b}$, $E^{l}_{r}$ and $E^{l}_{o}$ are causal resnet blocks for processing each condition, while $E^{l}_{s}$ is the shared causal resnet block.
Reference Conditions.
In urban scene generation, reference conditions $c^{ref}$ refer to the initial frame conditions which are optional and only used for future scene prediction similar to world models [29, 21]. We employ a 3D VAE encoder $E^{vae}$ [38] to encode these reference conditions but set the temporal dimension to 1 since each modality usually has only one initial frame. Note that in our approach, we employ a shared pre-trained 3D VAE from CogVideoX [38] for all modalities, rather than using different VAEs for different modalities. Since the 3D VAE in CogVideoX [38] is trained with a massive corpus of real-world videos, it performs well for encoding and decoding videos of different modalities. During model training, we use the shared pre-trained 3D VAE encoder to extract groundtruth frames latents to generate noisy latents. Therefore, the conditioned inputs $z$ is obtained by:
@@
Figure 3: An illustration of the proposed diffusion transformer block. Here, adaptive normalization, scaling and shifting are used between each layers to modulate latents.
@@ 3.3 Multi-Modal Multi-View Diffusion Transformer
Contemporary video diffusion models cannot be directly used for multi-view controllable urban scene generation. On the other hand, existing driving video diffusion models lack the ability to generate multi-modal data for holistic scene understanding. To resolve this problem, we devise a new multi-modal multi-view diffusion transformer model in this work.
Specifically, since a shared pre-trained 3D VAE $E^{vae}$ can be used to encode and decode videos of different modalities, we conjecture that different modalities share a common latent space and only require certain components to capture modal-specific content in order to distinguish them. In light of this, we decompose the model learning process into modal-shared learning and modal-specific learning and construct a unified diffusion transformer model with modal-shared components and modal-specific components.
For modal-shared components, we use temporal attention layers $D^{tem}$ with 3D full attention from CogVideoX [38] to learn temporal consistency across video frames and inject text conditions $f^{text}$ through cross-attention in the temporal attention layers. Here, the dimension of conditioned inputs $z$ of each modality is transformed from $\mathcal{R}^{V{\times}K{\times}H{\times}W{\times}C}$ to $\mathcal{R}^{V{\times}(NKW){\times}C}$ for the temporal attention layers. However, these temporal layers cannot guarantee spatiotemporal consistency for multi-view video generation. Hence, we append multi-view spatiotemporal blocks $D^{st}$ after temporal layers blocks to learn scene structure and capture spatiotemporal consistency. As shown in Fig. 3, each multi-view spatiotemporal block consists of a 3D spatial attention layer, a 3D spatial embedding layer, a spatiotemporal attention layer and a feedforward layer. The 3D spatial embedding layer is a multi-resolutional Hash grid [20] that encodes 3D occupancy locations $c^{occ}$ into 3D spatial embeddings, which are added to latents from the temporal layers block to enhance spatial consistency. The 3D spatial attention layer is a self-attention layer that transforms the latent dimension to $\mathcal{R}^{K{\times}(VHW){\times}C}$ to learn 3D spatial structure information of all surrounding camera views. The spatiotemporal attention layer is a 3D full attention layer that transforms the latent dimension to $\mathcal{R}^{(VKHW){\times}C}$ to capture full spatiotemporal information for multi-view driving scene generation. In practice, we append the multi-view spatiotemporal block after every $\alpha_{1}$ temporal layers block rather than every block. The feedforward layer is a fully-connected layer used to further transform latents. Note that following [38], we also employ adaptive normalization between each layer and adaptive scaling and shifting to modulate latents. We define this process as:
@@
For modal-specific components, we construct each cross-modal interaction layer block $D^{cm}_{m}$ with a self-attention layer, a cross-attention layer, and a feedforward layer, where adaptive normalization between each layer and scaling and shifting are also used. Here, the query of the cross-attention layer is the latents of each modality, while the key and the value are from the concatenated latents of the other modalities. This cross-modal interaction layer block is used to learn modal-specific content while maintaining cross-modal alignment. In practice, the cross-modal interaction layer block is appended after every $\alpha_{2}$ modal-shared layers. For each modality, this can process is formulated as:
@@
where $h^{\prime}_{m}$ is the modal-specific latents output from the cross-modal interaction layer, $h^{modal}_{m}$ denotes the key and the value from other modalities for the cross-attention layer. After $N$ modal-shared layers and cross-modal interaction layers, we employ the modal-specific projection head (a linear layer with adaptive normalization) to estimate the added noises $\epsilon$ and generate the clean latents $x^{\prime}$ for each modality. The clean latents $x^{\prime}$ for each modality is used as the input to a VAE decoder [38] to generate modal-specific multi-view scene videos. In this way, our diffusion transformer model is capable of generating multi-modal multi-view videos with high fidelity, controllability and spatiotemporal consistency, in a unified framework. Note that our approach can be used for both one-modal and multi-modal generation, and the modalities are not limited to RGB videos, depth maps, and semantic maps, but can be extended to other modalities, e.g., normal maps, flow maps, etc.

### v2

# MoVieDrive: Urban Scene Synthesis with Multi-Modal Multi-View Video Diffusion Transformer

@@ Abstract
Urban scene synthesis with video generation models has recently shown great potential for autonomous driving. Existing video generation approaches to autonomous driving primarily focus on RGB video generation and lack the ability to support multi-modal video generation. However, multi-modal data, such as depth maps and semantic maps, are crucial for holistic urban scene understanding in autonomous driving. Although it is feasible to use multiple models to generate different modalities, this increases the difficulty of model deployment and does not leverage complementary cues for multi-modal data generation. To address this problem, in this work, we propose a novel multi-modal multi-view video generation approach to autonomous driving. Specifically, we construct a unified diffusion transformer model composed of modal-shared components and modal-specific components. Then, we leverage diverse conditioning inputs to encode controllable scene structure and content cues into the multi-modal multi-view unified diffusion model. In this way, our approach is capable of generating multi-modal multi-view driving scene videos in a unified framework. Our thorough experiments on real-world autonomous driving dataset show that our approach achieves compelling video generation quality and controllability compared with state-of-the-art methods, while supporting multi-modal multi-view data generation.
Figure 1: An illustration of the proposed MoVieDrive approach to urban scene video generation in autonomous driving. Our approach can be used to generate multi-modal multi-view driving scene videos, to synthesize diverse driving scenes under diverse time and weather conditions, and to generate long videos without reference frames.
*David Huang contributed to this work during an internship at Huawei Canada.
@@ 1 Introduction
Urban scene video generation for autonomous driving has advanced rapidly in recent years. It can be used to generate controllable driving scenes, especially to synthesize long-tail scenarios that cannot be easily collected in the real world. This facilitates performance improvement and reliability evaluation in autonomous driving.
Contemporary video generation methods, such as SVD [2] and CogVideoX [40], have shown promising performance for generating high-quality videos. However, they cannot be directly used for autonomous driving because multi-view driving scene generation requires multi-view spatiotemporal consistency and high controllability. To address this problem, some recent studies have explored diffusion models [10, 25] for controllable multi-view urban scene generation and shown promising performance [31, 35, 7, 22]. However, these methods mostly focus on single-modal RGB video generation and lack the ability to support multi-modal video generation. As a perception-intensive task, autonomous driving has an inherent need for multi-modal data, such as depth maps and semantic maps, which can facilitate holistic urban scene understanding towards more efficient self-driving. A common approach to solving this problem is using multiple models to generate different modalities, but this often increases the difficulty of model deployment and does not fully exploit complementary cues for multi-modal data generation.
@@
In this work, we propose a novel multi-MOdal multi-VIEw video diffusion approach, dubbed MoVieDrive, to autonomous driving. Instead of using multiple models to generate different modalities, our approach proposes to employ a unified model for multi-modal multi-view driving scene generation. An illustration of our approach to urban scene video generation in autonomous driving is depicted in Fig. 1. Specifically, we employ diverse conditioning inputs, including text conditions, context reference conditions, and layout conditions, to guide controllable scene generation. Among them, text conditions guide holistic scene generation, context reference conditions are optional and only used for future scene prediction, and layout conditions capture fine-grained scene cues. Next, we decompose multi-modal multi-view scene generation into modal-shared learning and modal-specific learning, and devise a unified diffusion transformer model with modal-shared layers and modal-specific layers. This unified model alleviates the need for multiple models and improves model scalability. We then encode diverse conditions into the unified diffusion model for multi-modal multi-view scene generation. This helps to leverage complementary cues from multi-modal data for richer scene understanding to facilitate multi-modal multi-view urban scene generation in a unified framework. Experimental results on real-world autonomous driving dataset demonstrate that our approach can generate multi-modal multi-view urban scene videos with high fidelity and controllability, surpassing state-of-the-art methods. In summary, our contributions are:
  • We propose a novel framework that exploits diverse conditioning inputs and a diffusion transformer model for multi-modal multi-view autonomous driving scene generation. This fills a gap left by existing methods.
  • We introduce a novel unified multi-modal multi-view video diffusion transformer model composed of modal-shared components and modal-specific components model for urban scene video generation.
  • We provide thorough experiments on real-world autonomous driving dataset and demonstrate the superiority of our approach over state-of-the-art methods.
@@ 2 Related Work
The rapid advancement in urban scene and asset generation has significantly facilitated data synthesis and closed-loop evaluation in autonomous driving [39, 44, 8, 19]. There are two primary research directions in urban scene synthesis for autonomous driving. One line of work explores rendering methods, such as neural radiance fields and 3D Gaussian splatting, for urban scene synthesis [39, 44, 24, 36]. Despite high-fidelity scene synthesis, this type of method has poor controllability and diversity. Another line of work explores video generation, such as diffusion models [2, 40], for urban scene generation [35, 31]. Among them, some studies focus on single-view generation [9, 32], while others investigate multi-view generation [8, 33, 22]. Our work focuses on the more challenging multi-view setting. Contemporary methods, such as DriveDreamer [31], MagicDrive [8], MagicDrive-V2 [7], MaskGWM [22], etc., widely use diffusion models for multi-view urban scene generation. Although these methods have shown promising results, they only focus on RGB video generation and lack the ability to generate multi-modal data for holistic scene understanding. UniScene [14] proposes using multiple models to generate RGB videos and LiDAR point clouds, but still fails to construct a unified model for multi-view multi-modal urban scene generation. Unlike existing works, our approach proposes to generate multi-modal multi-view driving scene videos in a unified framework, which fills a gap left by existing methods.
@@
There have been several diffusion-based multi-modal synthesis methods proposed in recent years [18, 41, 27, 37]. However, none of them are designed for urban scene generation and modifying them for complex controllable multi-modal multi-view scene generation is not trivial. This often requires modifying conditioning input encoders, diffusion transformer layers, and training processes. Note that investigating how to modify these approaches to autonomous driving is beyond the scope of this work. Our approach differs from these works in that we propose a novel framework that uses diverse conditioning inputs to encode controllable scene structure and content cues into a diffusion transformer model for multi-modal multi-view driving scene generation.
@@ 3.1 Approach Overview
Fig. 2 shows an overview of our approach. Our approach follows the paradigm of diffusion-based video generation [7, 31]. As shown in the left of Fig. 2, we construct scene descriptions with text conditions, context reference conditions, and layout conditions, and employ the corresponding encoders to extract the embeddings of these conditions. The embeddings of layout conditions and context reference conditions are concatenated with noisy latents $x$ as inputs $z$ to a diffusion model, while the embeddings of text conditions are injected into the diffusion model through cross-attention layers. Here, noisy latents $x$ are obtained by sampling random noises from the Gaussian distribution during inference and by adding scheduled noises to the groundtruth frame latents extracted by a VAE encoder during training. Next, as shown in the middle of Fig. 2, we construct a multi-modal multi-view diffusion transformer model with modal-shared layers, composed of temporal layers and multi-view spatiotemporal blocks, and modal-specific layers, composed of cross-modal interaction layers and projection layers. We then train the diffusion transformer model to estimate the added noises $\epsilon{\sim}\mathcal{N}(0,I)$ from the conditioned inputs $z$, so as to generate clean latents $x^{\prime}$ for each modality. Finally, as shown in the right of Fig. 2, the clean latents $x^{\prime}$ are used as inputs to a VAE decoder to produce multi-view scene videos for each modality. In the following sections, we present the details of conditioning inputs encoding, multi-modal multi-view diffusion transformer, and model training and inference.
@@ 3.2 Conditioning Inputs Encoding
Following [7, 31], text conditions are used to guide holistic scene generation. In our framework, we employ two types of text conditions, including camera prompts and text prompts, and use two text encoders to generate text condition embeddings $f^{text}$. For camera prompts $c^{cam}$, we concatenate the camera intrinsic and extrinsic parameters applied with the Fourier embedding [20] and employ an MLP-based encoder $E^{cam}$ to extract camera embeddings. For text prompts $c^{text}$, we use video caption descriptions and apply a frozen text encoder $E^{text}$ to extract text embeddings. We concatenate these embeddings as $f^{text}$. This process can be formulated as:
@@
For fine-grained scene structure and content control, we employ three types of layout conditions, including box maps, road maps and occupancy-based layout maps. These layout conditions can be obtained from human annotations or pre-trained models and are useful to control fine-grained details in complex multi-view driving scene generation. For box maps, different from [8, 7] which use box coordinates to extract box embeddings, we directly generate 2D box maps $c^{b}$ by projecting target 3D box locations onto image planes and assign different colors to different categories. For road maps $c^{r}$, we project road structure cues (lane dividers, pedestrian crossings, and lane boundaries) onto image planes and assign different colors to different categories. In addition, following [15], we use occupancy-based layout maps $c^{o}$ to further enhance scene fine-grained control. This is accomplished by projecting sparse 3D occupancy onto an image plane to generate sparse semantic maps. Note that these semantic maps are coarse and sparse, which are different from the dense semantic modality used for urban scene understanding in autonomous driving. 3D occupancy has been well explored in autonomous driving, and in practice, 3D occupancy can be obtained from an off-the-shelf model [14] or a simulator [6].
@@
where $E^{l}_{b}$, $E^{l}_{r}$ and $E^{l}_{o}$ are causal resnet blocks for each condition, while $E^{l}_{s}$ is the shared causal resnet block.
Context Reference Conditions.
In urban scene generation, context reference conditions $c^{ref}$ refer to the initial frame conditions which are optional and only used for future scene prediction similar to world models [31, 22]. We employ a 3D VAE encoder $E^{vae}$ [40] to encode these context conditions but set the temporal dimension to one since each modality usually has only one initial frame for each camera view. In our approach, we employ a shared pre-trained 3D VAE from CogVideoX [40] for all modalities, rather than using different VAEs for different modalities. Since the 3D VAE in CogVideoX [40] is trained with a massive corpus of real-world videos, it performs well for encoding and decoding videos of different modalities. During model training, we use the shared pre-trained 3D VAE encoder to extract groundtruth frames latents to generate noisy latents. Therefore, the conditioned inputs $z$ is obtained by:
@@
Figure 3: An illustration of our diffusion transformer block.
@@ 3.3 Multi-modal Multi-view Diffusion Transformer
Contemporary video diffusion models cannot be directly used for multi-view controllable urban scene generation. On the other hand, existing driving video diffusion models lack the ability to generate multi-modal data for holistic scene understanding. To resolve this problem, we devise a new multi-modal multi-view diffusion transformer model. Specifically, since a shared pre-trained 3D VAE $E^{vae}$ can be used to encode and decode videos of different modalities, we conjecture that different modalities share a common latent space and only require certain components to capture modal-specific content in order to distinguish them. In light of this, we decompose the model learning process into modal-shared learning and modal-specific learning and construct a unified diffusion transformer model with modal-shared components and modal-specific components.
Modal-Shared Components.
We use temporal attention layers $D^{tem}$ with 3D full attention from CogVideoX [40] to learn temporal consistency across video frames and inject text conditions $f^{text}$ through cross-attention in the temporal attention layers. Here, the dimension of conditioned inputs $z$ of each modality is transformed from $\mathcal{R}^{V{\times}K{\times}H{\times}W{\times}C}$ to $\mathcal{R}^{V{\times}(NKW){\times}C}$ for the temporal attention layers. However, these temporal layers cannot guarantee spatiotemporal consistency for multi-view video generation. Hence, we append multi-view spatiotemporal blocks $D^{st}$ after temporal layers blocks to learn scene structure and capture spatiotemporal consistency. As shown in Fig. 3, each multi-view spatiotemporal block consists of a 3D spatial attention layer, a 3D spatial embedding layer, a spatiotemporal attention layer and a feedforward layer. The 3D spatial embedding layer is a multi-resolutional Hash grid [21] that encodes 3D occupancy locations $c^{occ}$ into 3D spatial embeddings, which are added to latents from the temporal layers block to enhance spatial consistency. The 3D spatial attention layer is a self-attention layer that transforms the latent dimension to $\mathcal{R}^{K{\times}(VHW){\times}C}$ to learn 3D spatial structure information of all surrounding camera views. The spatiotemporal attention layer is a 3D full attention layer that transforms the latent dimension to $\mathcal{R}^{(VKHW){\times}C}$ to capture full spatiotemporal information for multi-view driving scene generation. In practice, we append the multi-view spatiotemporal block after every $\alpha_{1}$ temporal layers block rather than every block. The feedforward layer is a fully-connected layer used to further transform latents. Following [40], we also employ adaptive normalization between each layer and adaptive scaling and shifting to modulate latents. We define this process as:
@@
Modal-Specific Components.
We construct each cross-modal interaction layer block $D^{cm}_{m}$ with a self-attention layer, a cross-attention layer, and a feedforward layer, where adaptive normalization between each layer and scaling and shifting are also used. Here, the query of the cross-attention layer is the latents of each modality, while the key and the value are from the concatenated latents of the other modalities. This cross-modal interaction layer block is used to learn modal-specific content while maintaining cross-modal alignment. In practice, the cross-modal interaction layer block is appended after every $\alpha_{2}$ modal-shared layers. This process for each modality is defined as:
@@
where $h^{\prime}_{m}$ is the modal-specific latents output from the cross-modal interaction layer, $h^{modal}_{m}$ denotes the key and the value from other modalities for the cross-attention layer. After $N$ modal-shared layers and cross-modal interaction layers, we employ the modal-specific projection head (a linear layer with adaptive normalization) to estimate the added noises $\epsilon$ and generate the clean latents $x^{\prime}$ for each modality. The clean latents $x^{\prime}$ for each modality is used as the input to a VAE decoder [40] to generate modal-specific multi-view scene videos. In this way, our diffusion transformer model can generate multi-modal multi-view videos with high controllability and spatiotemporal consistency in a unified framework. Our model can be used for both single-modal and multi-modal generation, and modalities are not limited to RGB videos, depth maps, and semantic maps.